### v1

# Nonreciprocal Transmission in Hybrid Atomic Ensemble-Optomechanical Systems

Abstract
We investigate perfect optical nonreciprocal transmission in a hybrid optomechanical system that incorporates an atomic ensemble. By introducing complex coupling strengths between the atomic ensemble and a mechanical oscillator, nonreciprocity is induced through interference between distinct optical pathways. The nonreciprocal transmission is governed by the real and imaginary components of the coupling constants, along with the relative phase differences between the optomechanical couplings. Our analysis reveals that, with precise tuning of system parameters, such as coupling strengths, detuning, and phase differences, perfect nonreciprocity can be achieved. We derive the conditions necessary for optimal nonreciprocal transmission and demonstrate its dependence on the complex nature of the coupling. These findings offer valuable insights for designing nonreciprocal optical devices, including isolators and circulators, with potential applications in quantum communication, signal processing, and photonics.
I Introduction
Interaction between atomic systems and mechanical oscillators has garnered considerable attention due to its potential in probing quantum phenomena and enabling novel applications. When coupled through radiation pressure, the mechanical system interacts with an optical field, constituting a cavity optomechanical (COM) system. These systems provide a versatile platform for exploring light-matter interactions at the quantum level. Notably, trapped atomic ensembles in optical cavities, which are strongly coupled to the optical field and mechanical oscillators, have demonstrated significant effects, such as quantum correlations and entanglement [1, 2, 3, 4, 5, 6]. These hybrid systems hold considerable promise for applications in quantum information processing [7, 8], mass sensing [9, 10, 11], fundamental investigations of quantum mechanics at macroscopic scales [12, 13], and nonlinear dynamics [14, 15].
A key functionality for quantum technologies is nonreciprocity, which facilitates unidirectional signal transmission and is pivotal in devices such as optical isolators and circulators. Such nonreciprocal devices are required for preventing undesired reflections and safeguarding sensitive optical components, such as lasers, from backscatter-induced instabilities [16, 17]. The realization of nonreciprocal transmission requires the breaking of time-reversal symmetry, a condition that has been theoretically and experimentally achieved through various mechanisms, including magnetic fields, optomechanical interactions, and engineered dissipative processes [18, 19, 20, 21, 22].
Recent advances in cavity optomechanics have unveiled new pathways for achieving nonreciprocal optical behavior without the reliance on magnetic effects. Specifically, robust optomechanical coupling has been demonstrated to induce nonreciprocal transparency and phase shifts in microring resonators [23]. Additionally, non-Hermitian physics has emerged as a potent framework for describing open quantum systems, wherein energy dissipation or gain plays a fundamental role. Non-Hermitian Hamiltonians, which incorporate complex coupling terms, have been extensively used to model nonreciprocal transmission in optical and mechanical systems, leveraging the interplay between loss and interference [24, 25, 20].
In this work, we investigate optical nonreciprocity in a hybrid atomic ensemble-cavity optomechanical system. By adjusting the system parameters—particularly the complex coupling strength $J_{3}$, which models dissipative interactions between the atomic ensemble and the mechanical mode—perfect nonreciprocity can be achieved. Our methodology employs a non-Hermitian Hamiltonian to describe the system, with its imaginary component $J_{3}$ symbolizing energy dissipation, resulting in unidirectional signal transmission. This is consistent with recent advances in non-Hermitian physics, where dissipation-induced nonreciprocity has been experimentally demonstrated in optical cavities and waveguides [20, 21].
The structure of this paper is as follows: In Section II, we introduce the theoretical model of the hybrid atomic ensemble-cavity optomechanical system and derive its Hamiltonian and equations of motion. In Section III, we analyze the conditions required for nonreciprocal transmission and study the effects of system parameters on the transmission properties. Finally, we conclude with a summary of our findings and discuss potential applications in quantum information processing and photonics.
II Perfect Optical Nonreciprocity
II.1 Model
We consider a hybrid optomechanical system consisting of two coupled cavities, one of which contains a trapped atomic ensemble, as illustrated in Figure 1. The first cavity mode is coupled to the atomic ensemble, while the two cavities are connected via photon tunneling. A mechanical oscillator is coupled to one of the cavities through radiation pressure, enabling interactions between the optical field and mechanical motion. The system is driven by strong external fields and probed by weak fields to analyze its optical response.
Figure 1: (a) A hybrid optomechanical system with atomic ensemble in one cavity. The coupling fields (probe fields) with amplitudes $\mathcal{E}_{1}$ and $\mathcal{E}_{2}~{}(\mathcal{E}_{p1}$ and $\mathcal{E}_{p2})$ are used to drive the two cavities that are coupled to each other via radiation pressure. (b) The diagram of interactions among subsystems in the optomechanical system.
The total Hamiltonian of the system include the following energy terms, the interactions among the subsystems, and the driving fields, and can be written as $(\hbar=1)$,
$H=H_{o}+H_{en}+H_{int}+H_{driv}.$ (1)
The term $H_{o}$ stands for the free Hamiltonian of the cavity fields and mechanical oscillator, and reads,
$H_{o}=\Delta_{1}a_{1}^{\dagger}a_{1}+\Delta_{2}a_{2}^{\dagger}a_{2}+\omega_{m}b^{\dagger}b,$ (2)
where $a_{j}(a_{j}^{\dagger})$ and $b(b^{\dagger})$ are the annihilation (creation) operators for the cavity and mechanical modes (with $j=1,2$), obeying the commutation relations $[a_{j},a_{j}^{\dagger}]=1$ and $[b,b^{\dagger}]=1$. The atomic ensemble, modeled as a bosonic mode, interacts with both the optical and mechanical subsystems. Its Hamiltonian is written as:
$H_{en}$ $=\frac{\Delta_{en}}{2}\sum_{j=1}^{N}\sigma_{j}^{z}+\left(J_{2}^{\ast}a_{1}\sum_{j=1}^{N}\sigma_{j}^{+}+J_{2}a_{1}^{\dagger}\sum_{j=1}^{N}\sigma_{j}^{-}\right)$ (3)
$+J_{3}\left(\sum_{j=1}^{N}\sigma_{j}^{+}+\sum_{j=1}^{N}\sigma_{j}^{-}\right)(b+b^{\dagger}),$
where $J_{2}$ is the coupling strength between the atomic ensemble and the first cavity mode, and $J_{3}$ describes the coupling between the atomic ensemble and the mechanical oscillator. The detuning of the atomic ensemble mode from the driving frequency is defined as $\Delta_{en}=\omega_{en}-\omega_{L}$. We assume that the number of atoms in the excited state is significantly lower than the total number of atoms $N$, indicating that the atomic ensemble operates under low excitation conditions. Consequently, we can treat the atomic ensemble as a bosonic mode. The atomic ensemble is initially modeled using spin-1/2 Pauli operators $\sigma_{j}^{z}$ and $\sigma_{j}^{\pm}$ for the $j^{th}$ two-level atom, which satisfy the commutation relations $[\sigma_{j}^{+},\sigma_{j}^{-}]=\sigma_{j}^{z}$ and $[\sigma_{j}^{z},\sigma_{j}^{\pm}]=\pm 2\sigma_{j}^{\pm}$. To treat the atomic ensemble as a bosonic mode, we apply the Holstein-Primakoff transformation, which enables us to map the collective spin operators to bosonic creation and annihilation operators,
$\frac{1}{N}\sum_{j=1}^{N}\sigma_{j}^{-}=d,$ $\frac{1}{N}\sum_{j=1}^{N}\sigma_{j}^{+}=d^{\dagger},$ $\sum_{j=1}^{N}\sigma_{j}^{z}=2d^{\dagger}d-N.$
Thus, the Hamiltonian for the atomic ensemble can be rewritten as:
$H_{en}=\Delta_{en}d^{\dagger}d+\left(J_{2}^{\ast}a_{1}d^{\dagger}+J_{2}a_{1}^{\dagger}d\right)+J_{3}(d+d^{\dagger})(b+b^{\dagger}).$ (4)
The interaction Hamiltonian between the cavity modes and the mechanical oscillator is given by,
$H_{int}=g_{1}a_{1}^{\dagger}a_{1}(b+b^{\dagger})+g_{2}a_{2}^{\dagger}a_{2}(b+b^{\dagger})+J_{1}(a_{1}a_{2}^{\dagger}+a_{2}a_{1}^{\dagger}),$ (5)
where $g_{j}(j=1,2)$ are the optomechanical coupling strengths between the cavity modes and the mechanical oscillator, and $J_{1}$ represents the photon tunneling coupling between the two cavities. The driving fields are modeled by the Hamiltonian,
$\begin{split}H_{driv}&=i(\mathcal{E}_{1}a_{1}^{\dagger}-\mathcal{E}_{1}^{\ast}a_{1})+i(\mathcal{E}_{2}a_{2}^{\dagger}-\mathcal{E}_{2}^{\ast}a_{2})\\ &+i\mathcal{E}_{p1}(a_{1}^{\dagger}e^{-i\delta t}-a_{1}e^{i\delta t})+i\mathcal{E}_{p2}(a_{2}^{\dagger}e^{-i\delta t}-a_{2}e^{i\delta t}),\end{split}$ (6)
where $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$ are the amplitudes of the driving fields for the cavities, $\mathcal{E}_{p1}$ and $\mathcal{E}_{p2}$ are the amplitudes of the probe fields, and $\delta$ is the detuning between the probe and driving fields.
The complex coupling strength $J_{3}$ plays a critical role in inducing nonreciprocal transmission. The real part of $J_{3}$ governs the interaction strength between the atomic ensemble and the mechanical oscillator, while the imaginary component represents dissipative processes, where energy is either lost to the environment or exchanged between subsystems. The introduction of dissipation through the imaginary part of $J_{3}$ breaks the symmetry of the system, leading to unidirectional signal transmission. This dissipation-induced nonreciprocity is a hallmark of non-Hermitian systems, which allows for nonreciprocal effects without the need for magnetic fields. For experimental realization of this non-Hermitian Hamiltonian, careful consideration must be given to coupling between atomic ensembles and mechanical modes. This can be achieved through engineered dissipative processes by tuning interaction strengths and decay rates in optomechanical systems with external baths.
II.2 Dynamical Equations
The dynamical evolution of the bosonic operators in the system is governed by the Heisenberg equations of motion, which are expressed as,
$\dot{a_{1}}$ $=-(i\Delta_{1}+\kappa_{1})a_{1}-ig_{1}a_{1}(b+b^{\dagger})-iJ_{1}a_{2}-iJ_{2}d+\mathcal{E}_{1}+\mathcal{E}_{p1}e^{-i\delta t},$ (7a)
$\dot{a_{2}}$ $=-(i\Delta_{2}+\kappa_{2})a_{2}-ig_{2}a_{2}(b+b^{\dagger})-iJ_{1}a_{1}+\mathcal{E}_{2}+\mathcal{E}_{p2}e^{-i\delta t},$ (7b)
$\dot{d}$ $=-(i\Delta_{en}+f)d-iJ^{\ast}_{2}a_{1}-iJ_{3}(b+b^{\dagger}),$ (7c)
$\dot{b}$ $=-(i\omega_{m}+\gamma)b-ig_{1}a_{1}^{\dagger}a_{1}-ig_{2}a_{2}^{\dagger}a_{2}-iJ_{3}(d+d^{\dagger}),$ (7d)
where the noise terms have been ignored. The first equation (Eq. (7a)) describes the dynamics of the optical field in cavity 1, which is influenced by the mechanical oscillator through radiation pressure (coupling strength $g_{1}$) and by photon tunneling to cavity 2 (with coupling strength $J_{1}$). Additionally, the atomic ensemble couples to cavity 1 with strength $J_{2}$, and external driving fields $\mathcal{E}_{1}$ and $\mathcal{E}_{p1}$ act on the cavity. The terms $i\Delta_{1}+\kappa_{1}$ account for the detuning and dissipation in cavity $1$. The second equation (Eq. (7b)) stands for the optical field in cavity $2$, and it follows a similar form but with distinct coupling strengths and driving terms, reflecting its interaction with cavity $1$, the mechanical mode, and the external field $\mathcal{E}_{2}$. The atomic ensemble mode (Eq. (7c)), represented by the operator $d$, evolves under the influence of the first cavity field $a_{1}$ (through coupling strength $J_{2}^{*}$) and the mechanical oscillator (via coupling strength $J_{3}$). The term $f$ models the dissipation within the atomic ensemble. The last equation (Eq. (7d)) governs the dynamics of the mechanical oscillator. Its motion is influenced by both optical fields (through strengths $g_{1}$ and $g_{2}$) and by its coupling to the atomic ensemble through $J_{3}$. The terms $i\omega_{m}+\gamma$ account for the mechanical frequency and damping.
To analyze the system dynamics, we use the standard linearization technique. Under strong optical driving conditions, where the steady-state values dominate and fluctuations are small, we expand each Heisenberg operator as the sum of its steady-state value and a small fluctuation around it:
$a_{j}=\alpha_{j}+\delta a_{j},$ $d=\rho+\delta d,$ $b=\beta+\delta b.$ (8)
Substituting these expansions into Eq. (7), we obtain two sets of equations: one for the steady-state mean values and another for the fluctuation operators, which govern the quantum dynamics. The steady-state mean values $\alpha_{j}$, $\rho$, and $\beta$ are found by solving the following system of equations:
$\begin{cases}(i\Delta^{\prime}_{1}+\kappa_{1})\alpha_{1}+iJ_{1}\alpha_{2}+iJ_{2}\rho-\mathcal{E}_{1}=0,\\ (i\Delta^{\prime}_{2}+\kappa_{2})\alpha_{2}+iJ_{1}\alpha_{1}-\mathcal{E}_{2}=0,\\ (i\Delta_{en}+f)\rho+iJ^{\ast}_{2}\alpha_{1}+2iJ_{3}\Re(\beta)=0,\\ (i\omega_{m}+\gamma)\beta+ig_{1}|\alpha_{1}|^{2}+ig_{2}|\alpha_{2}|^{2}+2iJ_{3}\Re(\rho)=0,\end{cases}$ (9)
where $\Delta^{\prime}_{j}=\Delta_{j}+2g_{j}\Re(\beta)$ are the effective detunings between the cavity modes and the coupling fields. In the steady-state regime, the mean values $\alpha_{j}$, $\rho$, and $\beta$ represent the amplitudes of the cavity fields, atomic ensemble, and mechanical displacement, respectively. The coupling terms between these subsystems introduce complex interactions: $J_{2}$ represents the interaction between the atomic ensemble and the cavity field, while $J_{3}$ models both conservative and dissipative interactions between the atomic ensemble and the mechanical oscillator. The steady-state equations indicate how energy is exchanged between these subsystems, with the imaginary part of $J_{3}$ driving dissipative energy exchange, which is critical for nonreciprocal behavior.
Next, we consider the quantum fluctuations around the steady state. The dynamics of the fluctuation operators are described by the linearized quantum Langevin equations, which take the form,
$\begin{cases}\delta\dot{a_{1}}=-(i\Delta^{\prime}_{1}+\kappa_{1})\delta a_{1}-i|J_{2}|e^{i\phi}\delta d-iJ_{1}\delta a_{2}\\ -iG_{1}(\delta b+\delta b^{\dagger})+\mathcal{E}_{p1}e^{-i\delta t},\\ \delta\dot{a_{2}}=-(i\Delta^{\prime}_{2}+\kappa_{2})\delta a_{2}-iJ_{1}\delta a_{1}-iG_{2}e^{i\theta}(\delta b+\delta b^{\dagger})\\ +\mathcal{E}_{p2}e^{-i\delta t},\\ \delta\dot{d}=-(i\Delta_{en}+f)\delta d-i|J_{2}|e^{-i\phi}\delta a_{1}-iJ_{3}(\delta b+\delta b^{\dagger}),\\ \delta\dot{b}=-(i\omega_{m}+\gamma)\delta b-iG_{1}(\delta a_{1}+\delta a^{\dagger}_{1})-iG_{2}(e^{-i\theta}\delta a_{2}\\ +e^{i\theta}\delta a^{\dagger}_{2})-iJ_{3}(\delta d+\delta d^{\dagger}),\end{cases}$ (10)
where the effective optomechanical coupling strengths are defined as $G_{1}=g_{1}\alpha_{1}$ and $G_{2}=g_{2}\alpha_{2}e^{-i\theta}$. The phase difference between the effective optomechanical couplings is $\theta=\theta_{2}-\theta_{1}$. Without loss of generality, we assume that $G_{1}$ is real, i.e., $\theta_{1}=0$, which reduces the phase difference to $\theta=\theta_{2}$. The quantum Langevin equations in Eq.(10) describe the small fluctuations around the steady-state values and govern the quantum dynamics of the system. These linearized equations incorporate the coupling between the optical, mechanical, and atomic subsystems, as well as dissipation via the terms involving $\kappa_{j}$, $\gamma$, and $f$. The coupling terms involving $J_{2}$ and $J_{3}$ play a central role in generating nonreciprocity, with the dissipation modeled by the imaginary part of $J_{3}$ leading to asymmetric transmission. The full quantum dynamics, including these fluctuations, can be used to calculate observables such as the transmission coefficients, which reveal the nonreciprocal behavior of the system. These calculations are essential for understanding how energy dissipation and interference in the system give rise to unidirectional optical transmission. To solve Eq. (10), we perform the transformations $\delta a_{j}\rightarrow\delta\tilde{a}_{j}e^{-i\Delta^{\prime}_{j}t}$, $\delta d\rightarrow\delta\tilde{d}e^{-i\Delta_{en}t}$, and $\delta b\rightarrow\delta\tilde{b}e^{-i\omega_{m}t}$. These transformations yield,
$\begin{cases}\delta\dot{\tilde{{a}}}_{1}=-\kappa_{1}\delta\tilde{a_{1}}-i|J_{2}|e^{i\phi}\delta\tilde{d}-iJ_{1}\delta\tilde{a_{2}}-iG_{1}\delta\tilde{b}+\mathcal{E}_{p1}e^{-iyt},\\ \delta\dot{\tilde{{a}}}_{2}=-\kappa_{2}\delta\tilde{a}_{2}-iJ_{1}\delta\tilde{a}_{1}-iG_{2}e^{i\theta}\delta\tilde{b}+\mathcal{E}_{p2}e^{-iyt},\\ \delta\dot{\tilde{d}}=-f\delta\tilde{d}-i|J_{2}|e^{-i\phi}\delta\tilde{a}_{1}-iJ_{3}\delta\tilde{b},\\ \delta\dot{\tilde{b}}=-\gamma\delta\tilde{b}-iG_{1}\delta\tilde{a}_{1}-iG_{2}e^{-i\theta}\delta\tilde{a}_{2}-iJ_{3}\delta\tilde{d}.\end{cases}$ (11)
We now assume the solution to Eq. (11) takes the form,
$\mathcal{O}=\mathcal{O}_{+}e^{-iyt}+\mathcal{O}_{-}e^{iyt}$ (12)
where $\mathcal{O}\equiv(\delta\tilde{a}_{j},\delta\tilde{d},\delta\tilde{b})$. Under the rotating wave approximation, where rapidly oscillating terms are neglected, we obtain a more compact matrix form as,
$\mathcal{A}_{1}\mathcal{X}=\mathcal{B},$ (13)
where,
$\mathcal{A}_{1}=\begin{pmatrix}\kappa_{1}-iy&iJ_{1}&i|J_{2}|e^{i\phi}&iG_{1}\\ iJ_{1}&\kappa_{2}-iy&0&iG_{2}e^{i\theta}\\ i|J_{2}|e^{-i\phi}&0&f-iy&iJ_{3}\\ iG_{1}&iG_{2}e^{-i\theta}&iJ_{3}&\gamma-iy\\ \end{pmatrix},$ (14)
$\mathcal{X}=\begin{pmatrix}\delta\tilde{a}_{1+}\\ \delta\tilde{a}_{2+}\\ \delta\tilde{d}_{+}\\ \delta\tilde{b}_{+}\\ \end{pmatrix},$ $\mathcal{B}=\begin{pmatrix}\mathcal{E}_{p1}\\ \mathcal{E}_{p2}\\ 0\\ 0\\ \end{pmatrix}.$ (15)
We can solve for $\mathcal{X}$ by inverting the matrix $\mathcal{A}_{1}$ as $\mathcal{X}=\mathcal{A}_{1}^{-1}\mathcal{B}$. This allows us to express the fluctuation components $\delta a_{1+}$ and $\delta a_{2+}$ as,
$\delta a_{1+}=\frac{(i\tau_{3}+\tau_{4})\mathcal{E}_{p1}+(i\tau_{1}-\tau_{2})\mathcal{E}_{p2}}{\mathcal{D}},$ (16)
$\delta a_{2+}=\frac{(i\chi_{1}-\chi_{2})\mathcal{E}_{p1}+(i\chi_{3}+\chi_{4})\mathcal{E}_{p2}}{\mathcal{D}},$ (17)
where
$\begin{cases}\tau_{1}=J_{1}y^{2}-J_{1}J^{2}_{3}-J_{1}\gamma f+G_{1}G_{2}ye^{-i\theta}+G_{2}|J_{2}|J_{3}e^{-i(\theta-\phi)}\\ \tau_{2}=J_{1}\gamma y+J_{1}fy+G_{1}G_{2}fe^{-i\theta},\\ \tau_{3}=-G^{2}_{2}y+y^{3}-J^{2}_{3}y-\gamma fy-\gamma y\kappa_{2}-fy\kappa_{2},\\ \tau_{4}=G^{2}_{2}f-\gamma y^{2}-fy^{2}-y^{2}\kappa_{2}+J^{2}_{3}\kappa_{2}+\gamma f\kappa_{2},\\ \chi_{1}=J_{1}y^{2}-J_{1}J^{2}_{3}-J_{1}\gamma f+G_{1}G_{2}ye^{i\theta}+G_{2}|J_{2}|J_{3}e^{i(\theta-\phi)},\\ \chi_{2}=J_{1}\gamma y+J_{1}fy+G_{1}G_{2}fe^{i\theta},\\ \chi_{3}=-G^{2}_{1}y+y^{3}-|J_{2}|^{2}y-G_{1}|J_{2}|J_{3}e^{i\phi}-J^{2}_{3}y-\gamma fy\\ -|J_{2}|G_{1}J_{3}e^{-i\phi}-\gamma y\kappa_{1}-fy\kappa_{1},\\ \chi_{4}=-\gamma y^{2}-fy^{2}-y^{2}\kappa_{1}+J^{2}_{3}\kappa_{1}+\gamma f\kappa_{1}+|J_{2}|^{2}\gamma+G^{2}_{1}f,\end{cases}$ (18)
and $\mathcal{D}$ is the determinant of $A_{1}$. Using the input-output formalism [26], the output field relations are given by:
$\mathcal{E}_{p1}^{out}+\mathcal{E}^{in}_{p1}e^{-iyt}=\sqrt{\kappa_{1}}\delta\tilde{a}_{1},$ (19)
$\mathcal{E}_{p2}^{out}+\mathcal{E}^{in}_{p2}e^{-iyt}=\sqrt{\kappa_{2}}\delta\tilde{a}_{2},$
where $\mathcal{E}_{p1,p2}^{in}=\mathcal{E}_{p1,p2}/\sqrt{\kappa_{j}}$. For simplicity, we assume that the output fields also take the form:
$\mathcal{O}=\mathcal{O}_{+}e^{-iyt}+\mathcal{O}_{-}e^{iyt}.$ (20)
By comparing the coefficients of the terms $e^{-iyt}$ and $e^{iyt}$, we obtain the following relations:
$\mathcal{E}^{out}_{p1+}=\sqrt{\kappa_{1}}\delta\tilde{a}_{1+}-\frac{\mathcal{E}_{p1}}{\sqrt{\kappa_{1}}},$ $\mathcal{E}_{p1-}^{out}=0,$ (21)
$\mathcal{E}^{out}_{p2+}=\sqrt{\kappa_{2}}\delta\tilde{a}_{2+}-\frac{\mathcal{E}_{p2}}{\sqrt{\kappa_{2}}},$ $\mathcal{E}_{p2-}^{out}=0.$
Particularly, perfect optical nonreciprocal transmission can be achieved if the transmission amplitudes satisfy the following conditions:
$\mathcal{T}_{12}=\left|\frac{\mathcal{E}_{p2+}^{out}}{\mathcal{E}_{p1}^{in}}\right|_{\mathcal{E}_{p2}^{in}=0}=1,$ $\mathcal{T}_{21}=\left|\frac{\mathcal{E}_{p1+}^{out}}{\mathcal{E}_{p2}^{in}}\right|_{\mathcal{E}_{p1}^{in}=0}=0,$ (22)
$\mathcal{T}_{12}=\left|\frac{\mathcal{E}_{p2+}^{out}}{\mathcal{E}_{p1}^{in}}\right|_{\mathcal{E}_{p2}^{in}=0}=0,$ $\mathcal{T}_{21}=\left|\frac{\mathcal{E}_{p1+}^{out}}{\mathcal{E}_{p2}^{in}}\right|_{\mathcal{E}_{p1}^{in}=0}=1.$
It is clear from these expressions that nonreciprocal transmission of optical signals can be achieved by adjusting the phase difference $\theta$ and the phase $\phi$ associated with the atoms. This will be explored in more detail in the next section.
III Results and Discussion
In this section, we analyze the conditions under which optical nonreciprocity can be achieved in the hybrid optomechanical system. We investigate the effects of key system parameters such as coupling strengths, detuning, dissipation, and phase differences on the transmission properties. The results reveal how tuning these parameters allows for control over nonreciprocal transmission, offering potential applications in optical isolation and signal routing.
III.1 Optical Nonreciprocal Transmission
The transmission amplitudes $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$ are defined to delineate the optical signal transmission from cavity 1 to cavity 2 and vice versa. These transmission coefficients are derived from the input-output relations and exhibit a strong dependence on the system parameters, particularly the phase differences between the effective optomechanical couplings and the coupling strengths of the atomic ensemble. Figure 2 displays the dependence of transmission amplitudes $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$ on the phase difference $\theta$ between the effective optomechanical couplings and the phase $\phi$ associated with the atomic ensemble. Notably, although the transmission amplitudes are not markedly influenced by phase $\phi$, both phases play a crucial role in the emergence of nonreciprocal transmission.
Figure 2: Density plot of transmission amplitudes ($\mathcal{T}_{12}$ and $\mathcal{T}_{21}$) versus relative phase difference $\theta$ between effective optomechanical coupling strength and the phase $\phi$. We have used $G_{1}/\gamma=0.5$, $G_{2}=G_{1}$, $f/\gamma=10$, $\kappa_{1}/\gamma=1$, $\kappa_{2}=\kappa_{1}$, $J_{1}/\gamma=0.5$, $\omega_{m}/\gamma=10$, $\Delta^{\prime}_{1}=\Delta^{\prime}_{2}=\Delta_{en}=\omega_{m}$, $|J_{2}|/\gamma=0.01$, and $J_{3}/\gamma=4.476e^{i\frac{\pi}{2}}$.
The results in Figure 2 suggest that optimal nonreciprocal transmission is realized when the coupling constant between the atomic ensemble and the mechanical oscillator, $J_{3}$, is complex. The complex nature of $J_{3}$ introduces dissipation, which is essential for breaking time-reversal symmetry and achieving nonreciprocity. Specifically, the imaginary component of $J_{3}$ encapsulates the energy exchange between the atomic ensemble and the mechanical oscillator, resulting in asymmetric optical signal transmission. From a physical standpoint, nonreciprocity arises from interference between two optical paths. One path is delineated along $a_{1}\rightarrow d\rightarrow b\rightarrow a_{2}$, while the other follows $a_{1}\rightarrow a_{2}$. The dissipative coupling $J_{3}$ enables unidirectional energy dissipation, thereby preventing back-propagation and facilitating nonreciprocal transmission.
III.2 Influence of Coupling Strength and Detuning
Figure 3: Transmission amplitudes $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$ as functions of detuning $y/\gamma$ for different effective phase differences $\theta$ and the phase $\phi$. (a) $\theta=\phi=0$, (b) $\theta=\phi=\frac{\pi}{4}$, (c) $\theta=\phi=\frac{\pi}{2}$, (d) $\theta=\phi=\frac{3\pi}{4}$, (e) $\theta=\phi=\pi$, (f) $\theta=\phi=\frac{5\pi}{4}$, (g) $\theta=\phi=\frac{3\pi}{2}$, and (h) $\theta=\phi=\frac{7\pi}{4}$. The other parameters are the same as in Figure 2.
We now examine how varying the coupling strengths $J_{2}$ and $J_{3}$, as well as the detuning $\Delta_{j}$, affect the transmission properties of the system. The study presented in Figure 3 shows the behavior of transmission amplitudes $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$ as functions of the detuning $y=\delta-\omega_{m}$ for different phase difference values $\theta$ and $\phi$. It is observed that when $\theta=\phi=m\pi$, where $m\in\mathbb{Z}$, the system exhibits reciprocal behavior, characterized by equal transmission amplitudes ($\mathcal{T}_{12}=\mathcal{T}_{21}$). This phenomenon occurs because the phase differences deplete the interference effect between the two optical paths. Conversely, when $\theta\neq m\pi$, the system manifests nonreciprocity with $\mathcal{T}_{12}\neq\mathcal{T}_{21}$. Optimal nonreciprocal transmission is achieved at resonance when $\theta=\phi=\frac{\pi}{2}$, where $\mathcal{T}_{12}\approx 0$ and $\mathcal{T}_{21}\approx 1$, or when $\theta=\phi=\frac{3\pi}{2}$, where $\mathcal{T}_{12}\approx 1$ and $\mathcal{T}_{21}\approx 0$. These findings illustrate that the direction of optical signal transmission can be regulated by adjusting the phase differences between the subsystems. Specifically, for $0<\theta<\pi$ and $0<\phi<\pi$, the observation of $\mathcal{T}_{12}<\mathcal{T}_{21}$ implies stronger forward transmission from cavity 2 to cavity 1. Conversely, for $\pi<\theta<2\pi$ and $\pi<\phi<2\pi$, the reverse holds, with $\mathcal{T}_{12}>\mathcal{T}_{21}$ indicating enhanced transmission in the opposite direction.
Figure 4: Transmission amplitudes $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$ as functions of detuning $y/\gamma$ for different coupling constants. (a) $|J_{2}|/\gamma=0.01$, (b) $|J_{2}|/\gamma=0.1$, (c) $|J_{2}|/\gamma=0.3$, (d) $|J_{2}|/\gamma=0.5$, (e) $J_{3}/\gamma=0.476e^{i\frac{\pi}{2}}$, (f) $J_{3}/\gamma=1.476e^{i\frac{\pi}{2}}$, (g) $J_{3}/\gamma=2.476e^{i\frac{\pi}{2}}$, and (h) $J_{3}/\gamma=4.476e^{i\frac{\pi}{2}}$. The other parameters are the same as in Figure 2, except for $\theta=\phi=\frac{\pi}{2}$.
Further investigation focuses on how variations in the coupling constants $|J_{2}|$ and $J_{3}$ affect the nonreciprocal response. As demonstrated in Figure 4, increasing the coupling constant $|J_{2}|$, which regulates the interaction between the atomic ensemble and the first cavity mode, primarily affects the transmission amplitude $\mathcal{T}_{12}$. Alternatively, decreasing the coupling constant $J_{3}$ between the atomic ensemble and the mechanical mode results in the reversal of input signal transmission, as illustrated in Figure 4(g-h).
These results underscore the critical reliance of nonreciprocity on the coupling constants. When $J_{3}$ is adjusted to possess a significant imaginary component, dissipation assumes a dominant role, inhibiting backscattered signals and facilitating strong nonreciprocal transmission. Conversely, reducing the dissipation (i.e., decreasing $\Im(J_{3})$) mitigates the nonreciprocal effect, prompting the system to display more reciprocal behavior.
III.3 Conditions for Perfect Nonreciprocity
We proceed to examine the required conditions for achieving perfect nonreciprocal transmission within the hybrid optomechanical system. Perfect nonreciprocity is realized when the transmission amplitude in one direction reaches unity ($\mathcal{T}_{12}=1$), while in the reverse direction it is completely suppressed ($\mathcal{T}_{21}=0$), or vice versa. This functionality is fundamental to applications in optical isolation and circulators within quantum information processing systems, where unidirectional light transmission is critical.
Perfect nonreciprocal transmission can be realized through strategic adjustments of several parameters: (i) the phase difference $\theta$ between the effective optomechanical couplings; (ii) the coupling strength $J_{3}$ between the atomic ensemble and the mechanical oscillator, with particular attention to its complex nature, where the real part signifies conservative interactions and the imaginary part indicates dissipation; (iii) the relative coupling strength $|J_{2}|$, which facilitates the interaction between the atomic ensemble and cavity 1; and (iv) the dissipation rates of the mechanical oscillator and the atomic ensemble, represented by $\gamma$ and $f$, respectively. The induction of sufficient dissipation through the imaginary part of $J_{3}$ is essential to disrupting time-reversal symmetry.
For simplicity, we consider the scenario where the phase differences are configured to $\theta=\phi=\frac{\pi}{2}$, a setup that generally yields pronounced nonreciprocal behavior. By defining $G_{1}=\sqrt{\gamma\kappa_{1}}$ and $G_{2}=\sqrt{\gamma\kappa_{2}}$, the necessary coupling strengths are determined by:
$\begin{cases}J_{1}=\frac{G_{1}G_{2}}{\gamma+f},\\ |J_{2}|=\frac{J_{1}J_{3}^{2}+J_{1}\gamma f-G_{1}G_{2}f}{G_{2}J_{3}},\\ J_{3}=\pm\sqrt{\frac{-R_{8}\pm\sqrt{R^{2}_{8}-4R_{7}R_{9}}}{2R_{7}}},\end{cases}$ (23)
The expressions above are obtained with the help of the conditions $\mathcal{T}_{12}=1$ and $\mathcal{T}_{21}=0$ (the expressions of $R_{7},R_{8},$ and $R_{9}$ are presented in Appendix). The dissipation introduced via the imaginary part of $J_{3}$ is instrumental in realizing perfect nonreciprocity. This dissipative interaction enables asymmetric energy exchange between the atomic ensemble and the mechanical oscillator, effectively mitigating back-reflections and facilitating unidirectional transmission. Conversely, reducing the dissipation (i.e., minimizing $\Im(J_{3})$) diminishes the nonreciprocal effect and tends to reinstate reciprocal transmission, where $\mathcal{T}_{12}=\mathcal{T}_{21}$.
Figure 5: Transmission amplitudes $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$ as functions of $y/\kappa_{2}$ for different values of atomic ensemble decay rate $f$. (a) $f/\kappa_{2}=0.1$, (b) $f/\kappa_{2}=1$, (c) $f/\kappa_{2}=5$. The other parameters are the same as in Figure 2, except for $\gamma/\kappa_{2}=0.01$, $\kappa_{1}/\kappa_{2}=10$, and $J_{3}=\sqrt{\frac{-R_{8}+\sqrt{R^{2}_{8}-4R_{7}R_{9}}}{2R_{7}}}$.
Figure 6: Transmission amplitudes $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$ as functions of $y/\kappa_{2}$ for different values of atomic ensemble decay rate $f$. (a) $f/\kappa_{2}=0.1$, (b) $f/\kappa_{2}=1$, (c) $f/\kappa_{2}=5$. The other parameters are the same as in Figure 2, except for $\gamma/\kappa_{2}=0.01$, $\kappa_{1}/\kappa_{2}=10$, and $J_{3}=-\sqrt{\frac{-R_{8}-\sqrt{R^{2}_{8}-4R_{7}R_{9}}}{2R_{7}}}$.
III.4 Influence of Atomic Ensemble Decay and Mechanical Damping
In order to further understand the impact of system parameters on nonreciprocal transmission, this study investigates the influence of the atomic ensemble decay rate $f$ and the mechanical damping rate $\gamma$ on the transmission amplitudes $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$. These parameters dictate the dissipation within the atomic and mechanical subsystems, respectively, and significantly affect the stability of nonreciprocity. Figures 5 and 6 illustrate that perfect nonreciprocity is maintained across a spectrum of atomic decay rates $f$, indicating that minor fluctuations in the decay rate do not substantially alter the transmission characteristics. Indeed, Figures 5(a) and 6(a) demonstrate that at reduced decay rates $f/\kappa_{2}=0.1$, the transmission amplitudes $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$ show pronounced peaks in proximity to resonance. Nonetheless, nonreciprocal transmission is confined to a limited detuning range, suggesting heightened sensitivity to slight detuning variations when the atomic decay rate is minimal. With an increase in the atomic ensemble decay rate to $f/\kappa_{2}=1$, as depicted in Figures 5(b) and 6(b), the range of detuning allowing for nonreciprocal transmission expands. This results in an amplification of transmission $\mathcal{T}_{12}$ while $\mathcal{T}_{21}$ decreases markedly, indicating enhanced nonreciprocity. This expanded detuning range suggests that the system is more resilient to changes in detuning. For elevated atomic decay rates ($f/\kappa_{2}=5$), as shown in Figures 5(c) and 6(c), the system approaches near-perfect nonreciprocity over a broader detuning range. The peak magnitudes of $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$ become less pronounced, implying that the system can maintain nonreciprocity. This infers that dissipation introduced by the atomic ensemble contributes a stabilizing effect in preserving nonreciprocal transmission. The findings from Figures 5 and 6 reveal that increased atomic ensemble decay rates $f$ yield wider nonreciprocal transmission bands, thereby strengthening the system’s resilience to detuning. This is consistent with the idea that increased dissipation helps suppressing back-reflections and fosters stronger nonreciprocity.
Figure 7: Transmission amplitudes $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$ as functions of detuning y$/\kappa_{2}$ for different values of mechanical damping rate $\gamma$. (a) $\gamma/\kappa_{2}=0.001$, (b) $\gamma/\kappa_{2}=0.01$, (c) $\gamma/\kappa_{2}=0.1$, (d) $\gamma/\kappa_{2}=1$. The other parameters are the same as in Figure 2, except for $f/\kappa_{2}=1$, $J_{3}=\sqrt{\frac{-R_{8}+\sqrt{R^{2}_{8}-4R_{7}R_{9}}}{2R_{7}}}$ and $\kappa_{1}/\kappa_{2}=10$.
Figure 8: Transmission amplitudes $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$ as functions of detuning y$/\kappa_{2}$ for different values of mechanical damping rate $\gamma$. (a) $\gamma/\kappa_{2}=0.001$, (b) $\gamma/\kappa_{2}=0.01$, (c) $\gamma/\kappa_{2}=0.1$, (d) $\gamma/\kappa_{2}=1$. The other parameters are the same as in Figure 2 except for $f/\kappa_{2}=1$, $J_{3}=-\sqrt{\frac{-R_{8}-\sqrt{R^{2}_{8}-4R_{7}R_{9}}}{2R_{7}}}$ and $\kappa_{1}/\kappa_{2}=10$.
Next, we investigate the influence of mechanical damping $\gamma$ on the transmission properties. Figures 7 and 8 present the transmission amplitudes $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$ as functions of detuning for different values of the mechanical damping rate $\gamma$. The mechanical damping rate $\gamma$ dictates the rate at which energy is dissipated in the mechanical oscillator. Upon varying $\gamma$, we observe the interesting behaviors in the transmission amplitudes. Indeed, for low mechanical damping rates ($\gamma/\kappa_{2}=0.001$, Figures 7(a) and 8(a)), the transmission amplitudes exhibit sharp peaks around resonance. The detuning range conducive to nonreciprocal transmission is narrow, rendering the system sensitive to minor variations in detuning. In this regime, mechanical damping is insufficient to substantially influence nonreciprocity. As the mechanical damping rate increases to $\gamma/\kappa_{2}=0.01$ (Figures 7(b) and 8(b)), there is a broadening of the nonreciprocal transmission range. The peaks of $\mathcal{T}_{12}$ become broader, and $\mathcal{T}_{21}$ decreases, indicating enhanced nonreciprocity. Mechanical damping assumes a more substantial role in modulating the system’s behavior, effectively suppressing back-propagating signals. At higher mechanical damping rates ($\gamma/\kappa_{2}=0.1$ and $\gamma/\kappa_{2}=1$, Figures 7(c-d) and 8(c-d)), the detuning range over which nonreciprocity occurs is significantly broadened. The peaks of $\mathcal{T}_{12}$ and $\mathcal{T}_{21}$ become flattened, suggesting that the system may uphold nonreciprocity over an extensive detuning range. In this regime, mechanical damping exerts sufficient influence to dominate the system’s dynamics, culminating in robust nonreciprocal transmission that is less susceptible to detuning variations.
The findings from Figures 7 and 8 illustrate that an increase in mechanical damping enhances the robustness of nonreciprocity by expanding the detuning range for nonreciprocal transmission. This suggests that mechanical dissipation serves as a stabilizing function, akin to atomic ensemble decay, in sustaining nonreciprocal transmission. Overall, the above results illustrate that through modulation of phase differences, coupling strengths, and dissipation rates, it is feasible to achieve and regulate perfect nonreciprocity in the hybrid optomechanical system. These outcomes provide a foundation for the design of nonreciprocal optical devices with implications in photonics and quantum information processing.
IV Conclusion
This paper presents a detailed analysis of a hybrid optomechanical system that incorporates an atomic ensemble to achieve perfect optical nonreciprocal transmission. By introducing complex coupling strengths between the atomic ensemble and the mechanical oscillator, it is demonstrated that nonreciprocal transmission results from interference among distinct optical paths within the system. Specifically, the study reveals that the real and imaginary components of the coupling constants, along with the relative phase differences among the optomechanical coupling strengths, are essential in determining the directionality of optical signals. Through meticulous tuning of system parameters—such as coupling strengths, detuning, and phase differences—optimal nonreciprocal transmission was achieved under specific resonance conditions. Furthermore, the conditions required for perfect nonreciprocity were derived, and explicit expressions for the critical system parameters were provided. These findings offer a foundational framework for designing optical isolators and circulators, which are essential components in quantum communication and signal processing applications. Future research may explore the experimental realization of this system, with a particular focus on the role of dissipative couplings and environmental interactions in maintaining nonreciprocity. Moreover, investigating the robustness of nonreciprocal transmission in the presence of system imperfections and noise would yield significant insights. Overall, the results advance the development of sophisticated nonreciprocal optical devices with prospective applications in quantum information processing, photonics, and precision metrology.
Acknowledgments
This work has been carried out under the Iso-Lomso Fellowship at Stellenbosch Institute for Advanced Study (STIAS), Wallenberg Research Centre at Stellenbosch University, Stellenbosch 7600, South Africa. K.S. Nisar is grateful to the funding from Prince Sattam bin Abdulaziz University, Saudi Arabia project number (PSAU/2024/R/1445). A.-H. A.-A. is thankful to the Deanship of Graduate Studies and Scientific Research at University of Bisha for supporting this work through the Fast-Track Research Support Program.
Author Contributions: K.B.E. and J.-X. conceptualized the work and carry out the simulations and analysis. A., P.D., and S.G.N.E. participated in all the discussions and provided useful suggestions to the final version of the manuscript. K.S.N. and A.-H.A.-A. supervised the work. All authors participated equally in the discussions and the preparation of the final manuscript.
Competing Interests: All authors declare no competing interests.
Data Availability
Relevant data are included in the manuscript and supporting information. Supplement data are available upon reasonable request.
Appendix: Expressions of $R_{7},R_{8}$, and $R_{9}$
The determinant of $A_{1}$ is given by the expression
$\mathcal{D}$ $=-2J_{1}|J_{2}|J_{3}G_{2}\cos(\theta-\phi)-2i|J_{2}|J_{3}\kappa_{2}G_{1}\cos\phi$ (24)
$-2iJ_{1}G_{1}G_{2}f\cos\theta-2|J_{2}|J_{3}G_{1}y\cos\phi-2G_{1}G_{2}y\cos\theta$
$+G^{2}_{2}\mathcal{D}_{1}+G^{2}_{1}\mathcal{D}_{2}+|J_{2}|^{2}\mathcal{D}_{3}+J^{2}_{1}\mathcal{D}_{4}+J^{2}_{3}\mathcal{D}_{5}+y^{3}\mathcal{D}_{6}$
$+y^{2}\mathcal{D}_{7}+\kappa_{1}\kappa_{2}\mathcal{D}_{8}-i\gamma fy\mathcal{D}_{9}+y^{4},$
where
$\mathcal{D}_{1}$ $=|J_{2}|^{2}-iy\kappa_{1}-ify+f\kappa_{1}-y^{2},$ (25)
$\mathcal{D}_{2}$ $=-y^{2}-ify-iy\kappa_{2}+f\kappa_{2},$
$\mathcal{D}_{3}$ $=-i\gamma y-iy\kappa_{2}+\gamma\kappa_{2}-y^{2},$
$\mathcal{D}_{4}$ $=-y^{2}+J^{2}_{3}-i\gamma y-ify+\gamma f,$
$\mathcal{D}_{5}$ $=-y^{2}-iy\kappa_{1}-iy\kappa_{2}+\kappa_{1}\kappa_{2},$
$\mathcal{D}_{6}$ $=i\kappa_{1}+i\kappa_{2}+i\gamma+if,$
$\mathcal{D}_{7}$ $=-\gamma f-\gamma\kappa_{2}-f\kappa_{1}-\kappa_{1}\kappa_{2}-f\kappa_{2}-\gamma\kappa_{1},$
$\mathcal{D}_{8}$ $=\gamma f-ify-i\gamma y,~{}~{}\mathcal{D}_{9}=\kappa_{1}+\kappa_{2}.$
For simplicity, we consider the case where $\theta=\phi=\frac{\pi}{2}$, and the expression of $\mathcal{D}$ becomes $\mathcal{D}=J_{3}^{2}(R_{5}+R_{2}^{\prime})+R_{4}+\frac{R_{6}}{J_{3}^{2}}(J_{1}J_{3}^{2}+J_{1}\gamma f-G_{1}G_{2}f)^{2}+R_{3}$, where $R_{1}=\frac{1}{\sqrt{\kappa_{1}\kappa_{2}}}$, $R_{2}=\frac{-2G_{1}G_{2}f}{R_{1}}$, $R^{\prime}_{2}=J_{1}^{2}+\kappa_{1}\kappa_{2}$, $R_{3}=G^{2}_{2}f\kappa_{1}+G_{1}^{2}f\kappa_{2}+J_{1}^{2}\gamma f+\kappa_{1}\kappa_{2}\gamma f$, $R_{4}=-2J_{1}f(J_{1}\gamma-G_{1}G_{2})$, $R_{5}=-2J_{1}^{2}$, $R_{6}=\frac{G_{2}^{2}+\gamma\kappa_{2}}{G_{2}^{2}}$, $R_{7}=R_{5}+R_{2}^{\prime}+R_{6}J_{1}^{2}$, $R_{8}=R_{4}-R_{2}+2R_{6}J_{1}^{2}\gamma f-2R_{6}J_{1}G_{1}G_{2}f+R_{3}$, and $R_{9}=R_{6}J_{1}^{2}\gamma^{2}f^{2}-2R_{6}J_{1}\gamma f^{2}G_{1}G_{2}+R_{6}G_{1}^{2}G_{2}^{2}f^{2}$.